# Obliviate: Neutralizing Task-agnostic Backdoors within the Parameter-efficient Fine-tuning Paradigm

### v1

@@ 1 Introduction
Recent works have explored the security implications of PEFT Hong and Wang (2023). For example, attackers inject backdoors into PLMs, and then activate the attacks on the final PEFT models. One of the most severe attacks on PEFT is task-agnostic backdoors, which manipulates the output representations of PLMs aiming to harm fine-tuned models for arbitrary downstream tasks. Shen et al. (2021); Chen et al. (2021a); Zhang et al. (2023); Du et al. (2023). This type of attack is less prone to forgetting backdoors when fine-tuning since PEFT freezes backdoored parameters of PLMs and updates only a minimal set of added parameters. Furthermore, the ability to adapt models to multiple downstream tasks magnifies the risk of task-agnostic backdoors.
@@ 2.2 Backdoor Attacks on PLMs
The backdoor attacks pose severe threats in the NLP domain, especially targeting LLMs Dai et al. (2019); Kurita et al. (2020); Chen et al. (2021b); Yan et al. (2023). Attackers compromise target models to misclassify the text inputs with textual triggers while properly working on the clean samples.
@@
More recently, the implications of backdoored PLMs on PEFT have raised concerns Hong and Wang (2023); Gu et al. (2023); Zhao et al. (2024). Notably, task-agnostic backdoor is particularly fatal for PEFT because: 1) PEFT freezes all the backdoored parameters of the PLMs, so that the PEFT models have difficulty in forgetting the backdoors via training the limited number of newly added parameters, 2) The primary role of PEFT is to efficiently adapt a PLM to diverse tasks. This poses a significant risk of task-agnostic backdoors, compromising multiple tasks by exploiting only a single backdoored model.
@@ 2.3 Backdoor Defenses
Poisoned sample detection. The traditional approach for backdoor defense is to detect poisoned samples that include triggers by observing their disparity with clean samples. STRIP Gao et al. (2021) determines poisoned samples based on the prediction entropy of perturbed inputs. RAP Yang et al. (2021) leverages the difference in prediction robustness between poisoned and clean samples. MDP Xi et al. (2023) applies a perturbation-based defense to few-shot prompt learning. PSIM Zhao et al. (2024) provides poisoned sample detection for LoRA and prompt-tuning. It rejects samples for which the model has high prediction confidence. Instead of entirely rejecting detected samples, ONION Qi et al. (2021a) removes the triggers from a given input by measuring its perplexity. However, these methods require large computation costs due to multiple predictions for each sample. Furthermore, implementing ONION and PSIM requires (task-specific) auxiliary models, which detracts from the advantages provided by PEFT.
Trigger inversion. The trigger inversion technique removes trigger-like embeddings from the inputs. In the NLP domain, existing methods Wang et al. (2019); Qiao et al. (2019); Tao et al. (2022); Xu et al. (2023) suffers from the discontinuity of sentences and the sparsity of embedding spaces. T-miner Azizi et al. (2021) is a sequence-to-sequence model for generating minimally transformed classifier inputs to induce misclassification. PICCOLO Liu et al. (2022b) addresses the discontinuity problem by changing the subject model to a differentiable form. DBS Shen et al. (2022) adopts a dynamically reducing temperature coefficient in the softmax function to make the optimizer focus the ground truth trigger. LMSanitator Wei et al. (2024) shows that existing trigger inversion methods are less effective in detecting task-agnostic backdoors. To address this problem, they invert the outputs of poisoned samples rather than inverting input triggers. However, it is limited to prompt-tuning schemes that train additional embeddings, which is not generally applicable to other PEFT architectures.
Model purification. Several researchers have made efforts to purify models to revert the misclassified results of poisoned samples. One simple solution is to fine-tune all the model parameters on sufficient clean samples, leveraging catastrophic forgetting of trigger information Shen et al. (2021). Neuron pruning is a more promising approach, which has been largely studied in the computer vision domain Liu et al. (2018); Wu and Wang (2021); Zeng et al. (2021). These methods refine backdoored models by removing or penalizing neurons related to backdoors. RECIPE Zhu et al. (2023) firstly adopts this idea to purify PLMs. Nevertheless, the neuron pruning approach is not suitable for PEFT; it directly modifies backdoored neurons of the PLMs that cannot be accessed by PEFT.
@@ 4.1 Design Intuition
Natural backdoor forgetting. Even though fine-tuning with clean samples is a fundamental defense strategy, PEFT shows challenges in forgetting backdoors effectively Hong and Wang (2023). To illustrate the differences between PEFT and full fine-tuning, we present an example of backdoored models in Figure 2. PEFT is limited to a small number of trainable parameters. Therefore, it struggles to eliminate the backdoors, resulting in an output that is still similar to the adversarial representation. In contrast, the fully fine-tuned model alters its outputs significantly, enabling correct prediction of the true label. The quantity of neurons trained on clean samples is important to separate model outputs from the adversarial representations.
@@ 4.2 Obliviate Details
Benign neuron amplification. Given the constraints on increasing trainable parameters in PEFT, we enhance the influence of neurons in PEFT layers to neutralize backdoors in PLMs. Our method is to amplify the magnitudes of these small yet benign parameters, relatively undermining the effectiveness of the PLM’s backdoored neurons. This is inspired by neuron amplification approaches Yu et al. (2023); Zhu et al. (2024), which involve scaling up neurons important to specific tasks (e.g., classification task on a clean dataset).
Figure 2: Outputs of models applying PEFT and full fine-tuning on backdoored PLMs.
@@
Therefore, we reduce the triggers’ attention scores through an optimization process, rather than eliminating them from the inputs. To this end, we introduce the attention regularization loss $\mathcal{L}_{reg}$ to decrease the $L_{2}$-norm of attention scores, thereby penalizing excessively high values among them:
@@
where $H_{i}$ denotes the set of attention heads in the $i$th transformer layer, $\textbf{a}_{h}$ represents the attention scores for each head, and the remaining notations are consistent with those used in Equation (1). Specifically, we focus on the attentions corresponding to certain output vectors. For sentence classification, we regularize the attention scores of input tokens on the [CLS] output. Although the training process involves only clean samples, this approach effectively reduces the influence of trigger tokens while preserving the original context information.
@@ 5.1.3 Metrics
Maximum ASR and average ASR. We additionally measure the maximum ASR (MASR) and average ASR (AASR) introduced by Zhu et al. (2023) to examine the best and overall attack performances that attackers can achieve when targeting a specific label, respectively.
@@ 5.1.4 Defense setup
Table 1: Defense performance against backdoors in RoBERTa models across PEFT architectures.
@@
5.1.5 Baselines
PSIM Zhao et al. (2024). PSIM identifies and rejects poisoned samples by focusing on those with abnormally high output confidences. We train the auxiliary model on each downstream task using the reset labels. We select the threshold by allowing a 2% drop in the CACC on the validation set.
@@ 5.2 Defense Performance
The experimental results for defending RoBERTa models against three backdoor attacks are illustrated in Table 1. Our defense method, Obliviate, effectively mitigates all the backdoors across various PEFT architectures, with the constraint of training only a minimal number of parameters. Especially, the LoRA layers account for just 0.47% of the the total parameters of RoBERTa. We achieve a considerable reduction in average ASR (83.6$\downarrow$) with only a minor impact on CACC (0.78$\downarrow$). Furthermore, our method shows significant reductions in MASR across all cases (93.3$\downarrow$), successfully neutralizing even the most effective triggers that can be selected by attackers. The defense is more effective in multiclass classification tasks such as AG News and HSOL than in SST-2, which is a binary classification task. We also verify the effectiveness of our defense method across natural language inference (NLI), named entity recognition (NER), and question and answering (QA) tasks, with detailed results illustrated in Appendix D. Additionally, the experimental results for BERT models are provided in Appendix E.
In comparison, the ONION approach demonstrates efficacy in mitigating task-agnostic backdoor attacks, especially on the AG News task. Nonetheless, it falls short of achieving the performance levels exhibited by our defense method. Unlike task-specific backdoors, which optimize predictions towards the target label, task-agnostic backdoors result in negligible variance in the output probabilities between clean and poisoned samples. Consequently, RAP fails to protect PEFT models from task-agnostic backdoors in most cases even though we permit a conservative FRR of 5% on clean samples. Similarly, PSIM leverages the confidence gap between clean and poisoned samples to address task-specific backdoors. Despite the careful selection of thresholds for PSIM, its defense capabilities remain unsatisfactory with few exceptions in cases of the POR attack.
Our method effectively counters task-agnostic backdoors that rely on pre-defined vectors and adversarial MLM. Furthermore, it shows great mitigation against the UOR attack, which optimizes adversarial outputs, as detailed in Appendix F. Our defense method dissociates model outputs from these optimized manipulations, demonstrating the effectiveness and versatility of our approach.
@@ 5.3 Output Representation Analysis
Table 2: Performance of PEFT models using benign PLMs on SST-2, with or without our defense.
@@
Table 3: Defense performance of RoBERTa model on SST-2 against task-specific backdoor attacks. We also present the results from using either the attention regularization loss (only reg) or the neuron amplification loss (only amp).
@@
5.4 Robustness of Defense Method
The defender considers that PLMs have potentially been compromised by task-agnostic backdoors. However, in real-world situations, defenders are often unaware of whether PLMs are backdoored or what types of attacks have been conducted. To demonstrate the robustness of our defense method, we assess its performance in practical scenarios.
5.4.1 Effects on benign PLMs
@@ 5.4.2 Defense against task-specific attacks
Defenders may develop PEFT models using PLMs that contain task-specific backdoors although these attacks are only effective when the attacker has knowledge of the downstream task. We evaluate the performance of our defense method against various task-specific backdoor attacks exploiting word triggers Hong and Wang (2023), syntactic structures Qi et al. (2021c), and style transfer Qi et al. (2021b), as shown in Table 3. Our method is particularly effective against the word-based backdoor attack, benefiting from both the benign neuron amplification and the attention score regularization techniques. Of the two techniques, the attention score regularization technique generally exhibits less significance in defending the syntactic and style backdoor attacks since it is specially designed to neutralize insertion-based triggers. Nevertheless, our method demonstrates moderate defense performance against both backdoors by amplifying the benign neurons within the PEFT layers. These results underscore the effectiveness and comprehensiveness of our approach.
@@ 5.4.3 Defense against adaptive attacks
Table 4: Defense performance of RoBERTa model on SST-2 against adaptive attacks.
@@ Appendix C Implementation Details
Defense setup. To adopt our defense method to PEFT, we follow the common training process of adapter Houlsby et al. (2019), LoRA Hu et al. (2021), and prefix-tuning Li and Liang (2021) as provided in their work. We utilize the PEFT implementations available in AdapterHub Pfeiffer et al. (2020). We use a batch size of 16 across all tasks. For the selection of $\lambda_{amp}$ and $\lambda_{reg}$ values, we select the highest values within a certain range that result in no more than a 2% drop in CACC on the validation set. The other hyperparameters are detailed in Table 6.
Table 6: Training hyperparameters for each PEFT architecture. % param: the proportion of trainable parameters in the RoBERTa models. Lr: learning rate.
@@
Baseline: PSIM. The w/o defense model of the baselines serves as the victim model. To train the defensive model for each downstream task, we create a dataset $\mathbb{D}_{\text{clean\_reset}}^{\text{train}}$ from the training set by resetting the labels. The proposed threshold $\gamma=0.7$ has shown to be mostly ineffective against task-agnostic backdoors. Therefore, we optimize it by selecting the smallest one from $\{0.52,0.55,0.6,0.62,0.65,0.7\}$, permitting a 2% drop in the CACC of the victim model on the validation set. If there is no threshold satisfying this criterion, we use the default value. For the multiclass classification tasks, we adjust the threshold to $\gamma/L*2$, where $L$ denotes the number of labels.
Appendix D Defense Performance on Additional Classification Tasks
Appendix E Defense Performance: BERT
Appendix F Defense against the UOR Attack
Table 7: Ablation study on RoBERTa adapter models for SST-2, without the neuron amplification loss (w/o amp) and the attention regularization loss (w/o reg).
@@ Appendix F Defense against the UOR Attack
Appendix G Ablation Study
@@
(a) Coefficient of the neuron amplification loss.
(b) Coefficient of the attention regularization loss.
Figure 8: Defense performance of adapter models using RoBERTa on SST-2 by adjusting defense loss coefficients.
Appendix H Impacts of Defense Loss Coefficients

### v3

@@ 1 Introduction
Recent works have explored the security implications of PEFT Hong and Wang (2023). For example, attackers can inject backdoors into PLMs, and then activate the attacks on the final PEFT models. One of the most severe attacks on PEFT is task-agnostic backdoors, which manipulates the output representations of PLMs aiming to harm fine-tuned models for arbitrary downstream tasks. Shen et al. (2021); Chen et al. (2021a); Zhang et al. (2023); Du et al. (2023). This attack is less prone to forgetting backdoors when fine-tuning since PEFT freezes backdoored PLM parameters and updates only a minimal set of added parameters. Furthermore, the ability to adapt models to multiple downstream tasks magnifies the risk of task-agnostic backdoors.
@@ 2.2 Backdoor Attacks on PLMs
The backdoor attacks pose severe threats in the NLP domain, especially targeting LLMs Dai et al. (2019); Kurita et al. (2020); Chen et al. (2021b); Yan et al. (2023); Song et al. (2024). Attackers compromise target models to misclassify the text inputs with textual triggers while properly working on the clean samples.
@@
More recently, the implications of backdoored PLMs on PEFT have raised concerns Hong and Wang (2023); Gu et al. (2023); Zhao et al. (2024). Notably, task-agnostic backdoor is particularly fatal for PEFT because: 1) PEFT freezes all the backdoored parameters of the PLMs, so that it has difficulty in forgetting the backdoors via training the limited number of added parameters, 2) The primary role of PEFT is to efficiently adapt a PLM to diverse tasks. This poses a significant risk of task-agnostic backdoors, compromising multiple tasks by exploiting only a single backdoored model.
@@ 2.3 Backdoor Defenses
The traditional approach for backdoor defense is to detect poisoned samples that include triggers by observing their disparity with clean samples, such as STRIP Gao et al. (2021), RAP Yang et al. (2021), and MDP Xi et al. (2023). PSIM Zhao et al. (2024) provides poisoned sample detection for LoRA and prompt-tuning. Instead of entirely rejecting detected samples, ONION Qi et al. (2021a) removes the triggers from the input by measuring its perplexity. However, these methods require large computation costs due to multiple predictions for each sample. Furthermore, ONION and PSIM requires (task-specific) auxiliary models, which undermines the advantages provided by PEFT.
Several researchers have made efforts to purify models to revert the misclassified results of poisoned samples. One simple solution is to fine-tune all the model parameters on sufficient clean samples, leveraging catastrophic forgetting of trigger information Shen et al. (2021). Neuron pruning is a more promising approach, which has been studied in the computer vision domain Liu et al. (2018); Wu and Wang (2021); Zeng et al. (2021). These methods refine backdoored models by penalizing neurons related to backdoors. However, they cause the PLM parameters to become task-specific, undermining the memory-efficient nature of PEFT. Purifying the PLM itself using a method like RECIPE Zhu et al. (2023) may be feasible. However, the negative impact of intensive purification on model performance is significant in the PEFT context.
@@ 4.1 Design Intuition
Natural backdoor forgetting. Even though fine-tuning with clean samples is a fundamental defense strategy, PEFT shows challenges in forgetting backdoors effectively Hong and Wang (2023). To illustrate the differences between PEFT and full fine-tuning, we present an example of RoBERTa models backdoored by the POR attack in Figure 2. PEFT is limited to a small number of trainable parameters. Therefore, it struggles to eliminate the backdoors, resulting in an output that is still similar to the adversarial representation. In contrast, the fully fine-tuned model alters its outputs significantly, enabling correct prediction of the true label. The quantity of neurons trained on clean samples is important to separate model outputs from the adversarial representations.
@@ 4.2 Obliviate Details
Benign neuron amplification. Given the constraints on increasing trainable parameters in PEFT, we enhance the influence of neurons in PEFT layers to neutralize backdoors in PLMs. Our method is to amplify the magnitudes of these small yet benign parameters, relatively undermining the effectiveness of the PLM’s backdoored neurons. This is inspired by neuron amplification techniques used for model merging Yu et al. (2023); Zhu et al. (2024).
@@
Figure 2: Outputs of models applying PEFT and full fine-tuning on backdoored PLMs in our experiments.
@@
Therefore, we reduce the triggers’ attention scores through an optimization process, rather than eliminating them from the inputs. To this end, we introduce the attention regularization loss $\mathcal{L}_{reg}$ to decrease the $L_{2}$-norm of attention scores, thereby penalizing excessively high values among them, expressed as follows:
@@
where $H_{i}$ denotes the set of attention heads in the $i$th transformer layer, $\textbf{a}_{h}$ represents the attention scores for each head, and the remaining notations are consistent with those used in Equation (1). Specifically, we focus on the attentions corresponding to certain output vectors. For sentence classification, we regularize the attention scores of input tokens on the [CLS] output. This approach allows the model to learn the context across all input tokens rather than focusing excessively on specific tokens. Consequently, although the training process involves only clean samples, this effect can be generalized to processing poisoned test samples.
@@ 5.1.3 Metrics
Maximum ASR and average ASR. We additionally measure the maximum ASR (MASR) and average ASR (AASR), as introduced in Zhu et al. (2023), to evaluate the best and overall attack performance that attackers can achieve when targeting a specific label.
@@ 5.1.4 Defense setup
5.1.5 Baselines
ONION Qi et al. (2021a). This defense method removes triggers from an input by identifying outlier words that reduce its perplexity.
RAP Yang et al. (2021). This backdoor defense leverages the robustness of prediction probabilities to identify poisoned samples.
PSIM Zhao et al. (2024). PSIM identifies and rejects poisoned samples by focusing on those with abnormally high output confidences.
Table 1: Defense performance against backdoors in the RoBERTa models across PEFT architectures.
@@ 5.2 Defense Performance
The experimental results for defending RoBERTa models against three backdoor attacks are illustrated in Table 1. Our defense method, Obliviate, effectively mitigates all the backdoors across the PEFT architectures, with the constraint of training only a minimal number of parameters. Especially, the LoRA layers account for just 0.47% of the the total parameters of RoBERTa. We achieve a considerable reduction in average ASR (83.6$\downarrow$) with only a minor impact on CACC (0.78$\downarrow$). Furthermore, our method shows significant reductions in MASR across all cases (93.3$\downarrow$), successfully neutralizing even the most effective triggers that can be selected by attackers. The defense is more effective in multiclass classification tasks such as AG News and HSOL than in SST-2, which is a binary classification task.
In comparison, ONION demonstrates efficacy in mitigating the task-agnostic backdoors, especially on the AG News task. Nonetheless, it falls short of achieving the performance levels exhibited by our defense method. Unlike task-specific backdoors, which optimize predictions towards the target label, task-agnostic backdoors result in negligible variance in the output probabilities between clean and poisoned samples. Consequently, RAP fails to protect the PEFT models from the task-agnostic backdoors in most cases even though we permit a conservative FRR of 5% on clean samples. Similarly, PSIM leverages the confidence gap between clean and poisoned samples. Despite the careful selection of thresholds, its defense capability remains unsatisfactory with few exceptions in cases of the POR attack. We provide further analysis of the training dynamics in Appendix D and the defense losses in Appendix E and F. We also verify the effectiveness of our defense method across natural language inference (NLI), named entity recognition (NER), and question and answering (QA) tasks, with detailed results illustrated in Appendix G. Additionally, the results for BERT models are provided in Appendix H.
Our method effectively counters task-agnostic backdoors that rely on pre-defined vectors and adversarial MLM. Furthermore, it shows great mitigation against the UOR attack, which optimizes adversarial outputs, as detailed in Appendix J. Our defense method dissociates model outputs from these optimized manipulations, demonstrating the effectiveness and versatility of our approach.
@@ 5.3 Output Representation Analysis
5.4 Robustness of Defense Method
In real-world situations, defenders are often unaware of whether PLMs are backdoored or what types of attack/defense have been conducted. To demonstrate the robustness of our defense method, we assess its performance in practical scenarios.
5.4.1 Effects on benign PLMs
Table 2: Performance of the PEFT models using benign PLMs on SST-2, with or without our defense.
@@ 5.4.1 Effects on benign PLMs
Table 3: Defense performance of the RoBERTa models using purified PLMs on SST-2 against the POR attack. Both refers to the results of subsequently applying our defense method to the purified PLMs.
PEFT Method CACC ASR MASR AASR
Adapter w/o def 92.26 100 100 99.94
RECIPE 90.72 32.38 51.85 13.37
Obliviate 91.10 5.18 2.96 2.26
Both 89.62 5.51 4.29 2.55
LoRA w/o def 93.30 100 100 95.06
RECIPE 89.68 25.05 39.77 12.21
Obliviate 91.16 9.04 9.07 5.29
Both 89.95 6.17 4.91 3.19
Prefix w/o def 92.26 100 100 98.94
RECIPE 89.79 12.72 11.22 5.56
Obliviate 91.21 4.64 3.12 2.39
Both 88.69 6.19 4.04 2.87
Table 4: Defense performance of the RoBERTa models on SST-2 against task-specific backdoors. We also present the results from using either the attention regularization loss (only reg) or the neuron amplification loss (only amp).
@@
5.4.2 Effects on purified PLMs
PLMs may have undergone purification using a method such as RECIPE Zhu et al. (2023). We first evaluate the defense performance of this method (see Table 3). The maximum number of purification epochs is set to allow for a 2% drop in CACC for the PEFT models. To assess the impact of purification, we also present the evaluation results of our defense method applied to the purified PLMs. RECIPE alone is insufficient for defending against task-agnostic backdoors. It requires undermining model performance to achieve strong backdoor removal, and this decline cannot be recovered during PEFT due to the highly regularized PLM parameters being frozen, unlike in full fine-tuning. Given this limitation of RECIPE in PEFT, applying both methods becomes redundant, as the defense effectiveness is primarily driven by our method.
@@ 5.4.3 Defense against task-specific attacks
Defenders may develop PEFT models using PLMs that contain task-specific backdoors although these attacks are only effective when the attacker has knowledge of the downstream task. We evaluate the performance of our defense method against various task-specific backdoors exploiting word triggers Hong and Wang (2023), syntactic structures Qi et al. (2021c), and style transfer Qi et al. (2021b), as shown in Table 4. Our method is particularly effective against the word-based backdoors, benefiting from both the benign neuron amplification and the attention score regularization techniques. Of the two techniques, the attention score regularization technique generally exhibits less significance in defending the syntactic and style backdoors since it is specially designed to neutralize insertion-based triggers. Nevertheless, our method demonstrates moderate defense performance against both backdoors by amplifying the benign neurons within the PEFT layers. These results underscore the effectiveness and comprehensiveness of our approach.
@@ 5.4.4 Defense against adaptive attacks
Table 5: Defense performance of the RoBERTa models on SST-2 against adaptive attacks.
@@ Appendix C Implementation Details
Table 7: Training hyperparameters for each PEFT architecture. % param: the proportion of trainable parameters in the RoBERTa models. Lr: learning rate. We exclude the reparameterization parameters of prefix-tuning from the proportion as they are removed after training.
@@
Defense setup. To adopt our defense method to PEFT, we follow the common training process of adapter Houlsby et al. (2019), LoRA Hu et al. (2021), and prefix-tuning Li and Liang (2021) as provided in their work. In the PEFT configuration, we set the rank $r$ to 16 for LoRA and the prefix length to 30 for prefix-tuning since setting either parameter too low leads to significant degradation in CACC, even without any defense. We utilize the PEFT implementations available in AdapterHub Pfeiffer et al. (2020). We use a batch size of 16 across all tasks. For the selection of $\lambda_{amp}$ and $\lambda_{reg}$ values, we select the highest values within a certain range that result in no more than a 2% drop in CACC on the validation set. The other hyperparameters are detailed in Table 7.
Baselines:
  • PSIM Zhao et al. (2024). PSIM identifies and rejects poisoned samples by focusing on those with abnormally high output confidences. The w/o defense model of the baselines serves as the victim model. To train the defensive model for each downstream task, we create a dataset $\mathbb{D}_{\text{clean\_reset}}^{\text{train}}$ from the training set by resetting the labels. The proposed threshold $\gamma=0.7$ has shown to be mostly ineffective against task-agnostic backdoors. Therefore, we optimize it by selecting the smallest one from $\{0.52,0.55,0.6,0.62,0.65,0.7\}$, permitting a 2% drop in the CACC of the victim model on the validation set. If there is no threshold satisfying this criterion, we use the default value. For the multiclass classification tasks, we adjust the threshold to $\gamma/L*2$, where $L$ denotes the number of labels.
(a) Adapter
(b) LoRA
(c) Prefix-tuning
Appendix D Training Dynamics
Table 8: Ablation study on RoBERTa the adapter models for SST-2, without the neuron amplification loss (w/o amp) and the attention regularization loss (w/o reg).
@@ Appendix D Training Dynamics
(a) Coefficient of the neuron amplification loss.
(b) Coefficient of the attention regularization loss.
Figure 9: Defense performance of the adapter models using RoBERTa on SST-2 by adjusting defense loss coefficients.
Appendix E Ablation Study
@@
Appendix F Impacts of Defense Loss Coefficients